# Data-driven Radiative Magnetohydrodynamics Simulations with the MURaM Code: the Emergence of Active Region 11158 and the X2.2 Flare

### v1

@@ Abstract
We present the application of the data-driven branch of the MURaM code to the extensively studied flare-productive active region 11158. We refine the hybrid model strategy, which was described in the earlier paper of this series, to model the emergence of the active region during 4 solar days starting shortly before 2011 February 11 and the eruption of an X2.2 flare on February 15. After 4 days of evolution, a major eruption of a magnetic flux rope occurs in the simulation at approximately 3 hours (3% difference) before the real flare. The eruption leads to magnetic reconnection that contributes to bulk heating in the chromosphere and corona. The deposition of flare energy in the chromosphere causes strong condensations and evaporations, which fill hot post-flare loops and bright flare ribbons that exhibit separation and extension similar to the observed ribbon evolution. The synthesized soft X-ray flux corresponds to X class, which is close to the real event. The upward eruption of the flux rope leads to a piston-driven shock and horizontal expansion that exerts a strong downward impact on the lower atmosphere and generate an apparently fast-propagating chromospheric Moreton wave. We conclude that the data-driven radiative simulation of this active region can reproduce the key observational results of the real flare and demonstrate the great potential of this method for studying solar eruptions in a realistic corona environment.
@@ 1 Introduction
Solar eruptions represent perhaps the most complex and dynamic interaction between solar magnetic field and plasma. The community has recognized the important application of numerical simulations that follow the observed evolution of photospheric magnetic fields and consider sophisticated physical processes for realistic plasma properties.(Patsourakos et al., 2020; Jiang et al., 2022; Guo et al., 2024; Jiang, 2024; Schmieder et al., 2024)
@@
Following previous papers in this series that described the implementation of the data-driven boundary in the Radiative MHD code MURaM (Chen et al., 2023a) and the application on non-eruptive emerging an active region (Chen, 2025), we apply this method to a well-known and extensively studied active region 11158, which produced the first X class flare in solar cycle 24 (Schrijver et al., 2011) and has been thoroughly analyzed from its emission properties (e.g., Young et al., 2013; Milligan et al., 2014) to magnetic structures (e.g., Liu & Schuck, 2012; Sun et al., 2012; Wang et al., 2012; Vemareddy et al., 2012; Tziotziou et al., 2013; Kazachenko et al., 2015; Sun et al., 2017).
@@
In particular, data-driven magnetohydrodynamic (MHD) simulations under different approximations have been employed to investigate the evolution and eruptions in this active region. Cheung & DeRosa (2012) and Lumme et al. (2017) conducted a magneto-frictional calculation to model the evolution of the coronal magnetic field over several days of emergence. These models omit the role of plasmas and represent a progress of methodology from a series of static magnetic field calculations (Jarolim et al., 2023) to a full MHD simulation that includes both magnetic fields and plasmas.
@@
Inoue et al. (2014, 2015) used a nonlinear-forcefree-field as the initial condition for an MHD simulation and modeled the magnetic evolution of during the X class flare, and Inoue et al. (2018) investigated the M6.6 flare that occurred two days earlier. Hayashi et al. (2018, 2019) conducted MHD simulations for approximately 15 hours with a focus on testing different methods of driving the boundary and the accumulation of magnetic energy in the coronal volume.
@@
Recently, MHD simulations with improved treatments for the plasma thermodynamics have become possible. The hybrid model of Afanasyev et al. (2023) covers the emergence of the magnetic field via a magneto-friction calculation and the eruption via an MHD simulation with thermal conduction and almost isothermal equation of state. Fan et al. (2024) improved the energy equation by including the missing radiative loss and more self-consistent heating to maintain the corona temperature. Their model setup has a sufficient spatial extent, which allows investigations of the large-scale impact of the eruption, such as coronal extreme-ultraviolet (EUV) waves. A global scale model of the X class flare in AR 11158 was reported by Jin et al. (2022), who investigated the large-scale dimming associated with coronal mass ejection (CME). Hoeksema et al. (2020) (and references therein) presented a comprehensive test of the capability of driving local/global scale magneto-frictional models and radiative MHD models that reproduced the pre-eruption state of AR 11158.
@@
The well-studied AR 11158 serves as an ideal example of testing and validating the data-driven branch of the MURaM code, which may provide new insight complementing the detailed studies that have been performed. This work is a part of a comprehensive project of an international collaboration to model evolving and erupting active regions. As the first attempt, we prefer to use a simple setup following Cheung & DeRosa (2012) to drive the evolution, which is sufficient to capture the most important magnetic structure for the eruption. The comprehensive model describes, at the same time, the eruption at the large-scale corona, as well as the detailed plasma dynamics in different layers of the atmosphere, which, to the best of our knowledge, has not been done in previous works. The rest of the paper is organized as follows. We described in Section 2 the multi-stage model strategy and setup of the numerical simulations. The results are presented in Section 3, and a summary and discussion are given in Section 4.
2 Methods
@@
We follow the method presented in our previous papers Chen et al. (2023a) and Chen (2025) to implement the data-driven bottom boundary and to conduct multi-stage simulations. Here, we describe the particular modification of the method and model setup for this active region.
2.1 The Data-Driven Boundary
@@
where $B_{z}$ is the radial component of the vector magnetic field observed by SDO/HMI (Scherrer et al., 2012; Hoeksema et al., 2014), $\Delta t=720$ s the cadence of the observations, and $\Omega=1.0\times 10^{-5}$ s-1. $\Omega$ represents a rotation applied on a vertical flux tube and is the primary parameter that controls the injection of free magnetic energy into the domain, as described in earlier papers (Cheung & DeRosa, 2012; Chen, 2025). We use an educated guess on the basis of the relative motion of the flux concentrations in this active region.
@@
The evolution of the vertical magnetic field in the photosphere is shown in the top row of Figure 1. This active region, as marked in Figure 1, comprises of two major sunspot pairs, P1–N1 and P2–N2  [^1] Sun et al. (2012) denoted them as P0–N0 and P1–N1., and a small parasitic bipole, P3–N3 (N3 merges with N2 and is not marked in the figure), that emerges after February 14, next to the northern edge of the N2 spot. The P3 flux concentration continues to intrude into the area surrounded by the N1, P2, and N2 spots, which causes the destruction of a pre-existing flux rope at the onset of the eruption, as demonstrated later in this paper. The relative shearing and collisional motion between two large sunspot pairs, particularly between N1 and P2, creates a strongly sheared polarity inversion line (PIL) and is suggested to be the key process for generating free magnetic energy in the corona (Toriumi et al., 2017; Chintzoglou et al., 2019; Rempel et al., 2023). We estimate the order of magnitude of $\Omega$ from the sunspot motion as follows. The shearing motion is treated as an equivalence of the N1 spot rotating around the P2 spot. In two days, N1 rotated by approximately $\pi/2$ rad with respect to P2, according to a comparison of Panels (b) and (d) in Figure 1 (see also Figure 3 in Sun et al., 2012), which corresponds to a rotation period of 8 days. In the case of solid body rotation, the vorticity of $1.0\times 10^{-5}$ s-1 corresponds to a period of 14.6 days. Control experiments with $\Omega=0$ or $6\times 10^{-6}$ s-1 yield no major eruptions by the end of February 15.
2.2 A Three-Stages Hybrid Strategy
@@
We modify the two-stage hybrid strategy described in Chen (2025) to address the combined evolution of the emergence of AR 11158, the formation of coronal plasma structures, and the onset and propagation of the eruption.
2.2.1 Zero-$\beta$ Model
@@
The computational domain of a zero-$\beta$ simulation is resolved by $256\times 256\times 1152$ grid points. The grid spacing are 576 km and 64 km in the horizontal and vertical directions, respectively. This setup represents a horizontal domain that is smaller than the actual active region by a factor of 1.875. As in the zero-$\beta$ run in Chen (2025), a speed up factor $f_{\rm sp}=12$ is applied to the time series of the observed magnetic field such that the interval of consecutive magnetograms becomes 60 s. The simulation starts from February 10, 22h UT and is evolved for approximately $6\times 10^{5}$ iterations, which covers an equivalent time period of almost 5 days. As shown in Section 3, a strong decay in the magnetic energy is detected at the equivalent time of approximately February 15, 0h UT. The properties of the eruption need to be modeled via sophisticated MHD simulations.
2.2.2 Radiative MHD Model for the Corona
@@
The second stage is a radiative MHD run (hereafter, evo. run) that models the coronal plasma before the eruption of interest. The speed up factor of time is NOT applied in the MHD simulations. The MHD run is started from approximately February 14, 20h UT on the basis of a zero-$\beta$ run snapshot.
@@
A major modification to the method described in Chen (2025) is that the MHD model represents the one-to-one size of the actual active region. The simulation domain is resolved by a $512\times 512\times 1920$ mesh, with a horizontal/vertical grid spacing of 540 km/{/}64 km. This setup leads to a domain width of 276.48 Mm and a height of 122.88 Mm. To construct the initial magnetic field, we calculate a potential field $\mathbf{B}^{\rm P}$ from the observed vertical magnetic field, with a periodic side boundary and the assumption of a vanishing field at infinity. It is convenient to calculate the non-potential field $\mathbf{B}^{\rm NP-256}$ from the zero-$\beta$ snapshot. The coordinates of this cube in all three direction are scaled up by a factor of 1.875 to match the size of the actual active region, and $\mathbf{B}^{\rm NP}$ is obtained by interpolating the non-potential field cube to the mesh of the MHD run. Finally the initial magnetic field is constructed by
$\mathbf{B}^{\rm init}=\mathbf{B}^{\rm P}+\mathbf{B}^{\rm NP}+\mathbf{B}^{\rm QS},$ (3)
@@
where $\mathbf{B}^{\rm QS}$ is the magnetic field adapted from a fully relaxed quiet Sun simulation similar to that shown in Chen et al. (2022).
@@
The initial plasma conditions are also adapted from the quiet Sun simulation. As we found in simulations for a non-erupting active region (Chen, 2025), the active region corona evolves on the basis of the heating in the coronal volume and is almost independent of the initial plasma conditions.
@@
The MHD model is evolved untill the expected eruption of interest occurs, which takes approximately 3.2 hours. We set up a new MHD model with a further extend vertical domain to model the dynamics of the eruption.
2.2.3 Radiative MHD Model for the Eruption
@@
The third stage in the hybrid model is an MHD simulation with an extend vertical domain (hereafter, flare run). The flare run has a mesh of $512\times 512\times 3840$ grid points with the same grid spacing as the evo. run; thus, the domain height is extended to 245.76 Mm.
@@
3 Results
3.1 Emergence of the Active Region and Energy Budget of the Erution
@@
The coronal magnetic field in the zero-$\beta$ run across 4 days [^2] The first three columns show the snapshots corresponding to about $0^{\rm h}$ UT of each day, whereas the last column shows the snapshot corresponding to February 14, $20^{\rm h}$ UT, which is identical to the initial state of the MHD evo. simulation. are displayed in Panels (e) – (h) of Figure 1. The seed points for calculating fieldlines are uniformly distributed in the rectangular area covering the sunspots in this active region. In the latter two days, the coronal fieldlines closely resemble the shape of observed EUV loops. However, a quantitative investigation of the formation of these loops requires a model with much finer resolution, as suggested in our earlier study on a non-eruptive active region (Chen, 2025).
@@
A crucial parameter of flare-productive active regions is the free magnetic energy stored in the corona. Panel (i) of Figure 1 plots the evolution of volume-integrated total magnetic energy $E_{\rm tot}$ and free magnetic energy $E_{\rm free}$ in the zero-$\beta$ run [^3] On the basis of the setup of the zero-$\beta$ run, the magnetic energy and time are multiplied by a factor of 1.8753 and $12$, respectively, such that they can be directly compared with quantities of the MHD simulations. The rapid increase in the magnetic energy starts from February 13, similar to the trend found in observations (Sun et al., 2012). In our model, the magnetic energy increases to approximately $1.4\times 10^{33}$ erg with a very steady rate untill a time corresponding to February 14, $23^{\rm h}$ UT; meanwhile, the free magnetic energy accumulates to approximately $6.5\times 10^{32}$ erg. We note that this energy includes a large contribution from the curved/twist magnetic field outside the core region of the domain, which is an effect of the constant $\Omega$ and periodic side boundary. An abrupt decay in the free magnetic energy, as indicated by the gray mask in Panel (i), indicates a major eruption, which occurs approximately 3 hours earlier than the actual X2.2 flare on real Sun.
@@
The self-consistent evolution of the plasma and magnetic fields from development to eruption is described by the evo. and flare runs. Figure 1(j) shows the joint evolution of the total and free magnetic energy since 2.5 hours before the flare onset. The time $t$ hereafter are relative to the beginning of the MHD flare run. The free magnetic energy steadily increases and starts to decrease at $t=219$ s. The fast decay lasts for approximately 1000 s, during which $1.94\times 10^{32}$ erg magnetic energy is released.
3.2 Magnetic Structure and Plasma Dynamics of the Eruption

### v2

@@ Abstract
We present the application of the data-driven branch of the MURaM code to the extensively studied flare-productive active region 11158. We refine the hybrid model strategy, which was described in the earlier paper of this series, to model the emergence of the active region during 4 solar days starting shortly before 2011 February 11 and the eruption of an X2.2 flare on February 15. After 4 days of evolution, a major eruption of a magnetic flux rope occurs in the simulation at approximately 3 hours (3% difference) before the real flare. The eruption leads to magnetic reconnection, which contributes to bulk heating in the chromosphere and corona. The deposition of flare energy in the chromosphere causes strong condensations and evaporations, which fill hot post-flare loops and bright flare ribbons that exhibit separation and extension similar to the observed ribbon evolution. The synthesized soft X-ray flux corresponds to X class, which is close to the real event. The upward eruption of the flux rope leads to a piston-driven shock and horizontal expansion that exerts a strong downward impact on the lower atmosphere and generate an apparently fast-propagating chromospheric Moreton wave. We conclude that the data-driven radiative simulation of this active region can reproduce the key observational results of the real flare and demonstrate the great potential of this method for studying solar eruptions in a realistic corona environment.
@@ I Introduction
Solar eruptions represent perhaps the most complex and dynamic interaction between the solar magnetic field and plasma. The community has recognized the important application of numerical simulations that follow the observed evolution of photospheric magnetic fields and consider sophisticated physical processes for realistic plasma properties.(S. Patsourakos et al., 2020; C. Jiang et al., 2022; Y. Guo et al., 2024; C. Jiang, 2024; B. Schmieder et al., 2024)
@@
Following previous papers in this series that described the implementation of the data-driven boundary in the Radiative MHD code MURaM (F. Chen et al., 2023a) and the application on non-eruptive emerging an active region (F. Chen, 2025), we apply this method to a well-known and extensively studied active region 11158, which produced the first X class flare in solar cycle 24 (C. J. Schrijver et al., 2011) and has been thoroughly analyzed from its emission properties (e.g., P. R. Young et al., 2013; R. O. Milligan et al., 2014) to magnetic structures (e.g., Y. Liu & P. W. Schuck, 2012; X. Sun et al., 2012; S. Wang et al., 2012; P. Vemareddy et al., 2012; K. Tziotziou et al., 2013; M. D. Kazachenko et al., 2015; X. Sun et al., 2017).
@@
In particular, data-driven magnetohydrodynamic (MHD) simulations under different approximations have been employed to investigate the evolution of and eruptions in this active region. M. C. M. Cheung & M. L. DeRosa (2012) and E. Lumme et al. (2017) conducted a magneto-frictional calculation to model the evolution of the coronal magnetic field over several days of emergence. These models omit the role of plasmas and represent the progress of methodology from a series of static magnetic field calculations (R. Jarolim et al., 2023) to a full MHD simulation that includes both magnetic fields and plasmas.
@@
S. Inoue et al. (2014, 2015) used a nonlinear-forcefree-field as the initial condition for an MHD simulation and modeled the magnetic evolution during the X class flare, and S. Inoue et al. (2018) investigated the M6.6 flare that occurred two days earlier. K. Hayashi et al. (2018, 2019) conducted MHD simulations for approximately 15 hours with a focus on testing different methods of driving the boundary and the accumulation of magnetic energy in the coronal volume.
@@
Recently, MHD simulations with improved treatments for plasma thermodynamics have become possible. The hybrid model of A. N. Afanasyev et al. (2023) covers the emergence of the magnetic field via a data-driven magneto-friction calculation and the eruption via a data-constrained MHD simulation with thermal conduction and an almost isothermal equation of state. Y. Fan et al. (2024) improved the energy equation by including the missing radiative loss and more self-consistent heating to maintain the corona temperature. Their model setup has a sufficient spatial extent, which allows investigations of the large-scale impact of the eruption, such as coronal extreme-ultraviolet (EUV) waves. A global scale model of the X class flare in AR 11158 was reported by M. Jin et al. (2022), who investigated the large-scale dimming associated with coronal mass ejection (CME). J. T. Hoeksema et al. (2020) (and references therein) presented a comprehensive test of the capability of driving local/global scale magneto-frictional models and radiative MHD models that reproduced the pre-eruption state of AR 11158.
@@
The well-studied AR 11158 serves as an ideal example of testing and validating the data-driven branch of the MURaM code, which may provide new insight complementing the detailed studies that have been performed. This work is a part of a comprehensive project of an international collaboration to model evolving and erupting active regions. As the first attempt, we prefer to use a simple setup following M. C. M. Cheung & M. L. DeRosa (2012) to drive the evolution, which is sufficient to capture the most important magnetic structure for the eruption. The comprehensive model describes, at the same time, the eruption at the large-scale corona, as well as the detailed plasma dynamics in different layers of the atmosphere, which, to the best of our knowledge, has not been done in previous works. The rest of the paper is organized as follows. We described in Section II the multi-stage model strategy and setup of the numerical simulations. The results are presented in Section III, and a summary and discussion are given in Section IV.
II Methods
@@
We follow the method presented in our previous papers F. Chen et al. (2023a) and F. Chen (2025) to implement the data-driven bottom boundary and to conduct multi-stage simulations. Here, we describe the particular modification of the method and model setup for this active region.
II.1 The Data-Driven Boundary
@@
where $B_{z}$ is the radial component of the vector magnetic field observed by SDO/HMI (P. H. Scherrer et al., 2012; J. T. Hoeksema et al., 2014), $\Delta t=720$ s the cadence of the observations, and $\Omega=1.0\times 10^{-5}$ s-1. $\Omega$ represents a rotation applied on a vertical flux tube and is the primary parameter that controls the injection of free magnetic energy into the domain, as described in earlier papers (M. C. M. Cheung & M. L. DeRosa, 2012; F. Chen, 2025). We use an educated guess on the basis of the relative motion of the flux concentrations in this active region to determine the value of $\Omega$ as follows.
@@
The evolution of the vertical magnetic field in the photosphere is shown in the top row of Figure 1. This active region, as marked in Figure 1, comprises two major sunspot pairs, P1–N1 and P2–N2  [^1] X. Sun et al. (2012) denoted them as P0–N0 and P1–N1., and a small parasitic bipole, P3–N3 (N3 merges with N2 and is not marked in the figure), that emerges after February 14, next to the northern edge of the N2 spot. The P3 flux concentration continues to intrude into the area surrounded by the N1, P2, and N2 spots, which causes the destruction of a pre-existing flux rope at the onset of the eruption, as demonstrated later in this paper. The relative shearing and collisional motion between two large sunspot pairs, particularly between N1 and P2, creates a strongly sheared polarity inversion line (PIL) and is suggested to be the key process for generating free magnetic energy in the corona (S. Toriumi et al., 2017; G. Chintzoglou et al., 2019; M. Rempel et al., 2023). We estimate the order of magnitude of $\Omega$ from the sunspot motion as follows. The shearing motion is treated as an equivalence of the N1 spot rotating around the P2 spot. In two days, N1 rotated by approximately $\pi/2$ rad with respect to P2, according to a comparison of Panels (b) and (d) in Figure 1 (see also Figure 3 in X. Sun et al., 2012), which corresponds to a rotation period of 8 days. In the case of solid body rotation, the vorticity of $1.0\times 10^{-5}$ s-1 corresponds to a period of 14.6 days. Control experiments with $\Omega=0$ or $6\times 10^{-6}$ s-1 yield no major eruptions by the end of February 15.
II.2 A Three-Stages Hybrid Strategy
@@
We modify the two-stage hybrid strategy described in F. Chen (2025) to address the combined evolution of the emergence of AR 11158, the formation of coronal plasma structures, and the onset and propagation of the eruption.
II.2.1 Zero-$\beta$ Model
@@
The computational domain of a zero-$\beta$ simulation is resolved by $256\times 256\times 1152$ grid points. The grid spacings are 576 km and 64 km in the horizontal and vertical directions, respectively. This setup represents a horizontal domain that is smaller than the actual active region by a factor of 1.875. As in the zero-$\beta$ run in F. Chen (2025), a speed-up factor $f_{\rm sp}=12$ is applied to the time series of the observed magnetic field such that the interval of consecutive magnetograms becomes 60 s. This option is motivated by the fact that the evolution of the coronal magnetic field is much faster than the photosphere flux emergence and provides a significant reduction in computation time. Moreover, control experiments that employ smaller $f_{\rm sp}$ and the actual active region size yield consistent results in the magnetic energy evolution and the onset of a major eruption, as discussed in Appendix A. The simulation starts from February 10, 22h UT and is evolved for approximately $6\times 10^{5}$ iterations, which covers an equivalent time period of almost 5 days. As shown in Section III, a strong decay in the magnetic energy is detected at the equivalent time of approximately February 15, 0h UT. The eruption, which involves a consistent evolution of the plasma and magnetic field, needs to be modeled via MHD simulations as described below.
II.2.2 Radiative MHD Model for the Corona
@@
The second stage is a radiative MHD run (hereafter, evo. run) that models the coronal plasma before the eruption of interest. The speed-up factor of time is NOT applied in the MHD simulations. The MHD run is started from approximately February 14, 20h UT on the basis of a zero-$\beta$ run snapshot.
@@
A major modification to the method described in F. Chen (2025) is that the MHD model represents the one-to-one size of the actual active region. The simulation domain is resolved by a $512\times 512\times 1920$ mesh, with a horizontal/vertical grid spacing of 540 km/{/}64 km. This setup leads to a domain width of 276.48 Mm and a height of 122.88 Mm. To construct the initial magnetic field, we calculate a potential field $\mathbf{B}_{\rm P-512}$ from the observed vertical magnetic field, with a periodic side boundary and the assumption of a vanishing field at infinity. The non-potential field of the zero-$\beta$ snapshot $\mathbf{B}_{\rm NP-256}$ can be obtained by deducting the potential field component $\mathbf{B}_{\rm P-256}$ from zero-$\beta$ magnetic field cubes $\mathbf{B}_{\rm z\beta}$. The coordinates of $\mathbf{B}_{\rm NP-256}$ in all three directions are scaled up by a factor of 1.875 to match the size of the actual active region. Then, the cube is interpolated to a $512\times 512\times 2160$ mesh to fit the grid spacings of the MHD run, and we extract the lower 1920 grid points as $\mathbf{B}_{\rm NP-512}$. Finally, the initial magnetic field is constructed by
$\mathbf{B}_{\rm init}=\mathbf{B}_{\rm P-512}+\mathbf{B}_{\rm NP-512}+\mathbf{B}_{\rm QS},$ (3)
@@
where $\mathbf{B}_{\rm QS}$ is the magnetic field adapted from the part of the domain starting from the photosphere in a fully relaxed quiet Sun simulation similar to that shown in F. Chen et al. (2022). More specifically, the lower 1800 grid points of the $\mathbf{B}_{\rm QS}$ cube are extracted from the quiet Sun snapshot, and the upper 120 grid points are filled with a potential field that is calculated from the vertical magnetic field of the top layer.
@@
The initial plasma conditions, including the density, internal energy, and velocities, are also adapted from the quiet Sun simulation in a similar fashion. The thermodynamic variable of upper 120 grid points is filled by a hydrostatic equilibrium and the velocities are set as zero. As we found in simulations for a non-erupting active region (F. Chen, 2025), the active region corona evolves on the basis of the heating in the coronal volume and is almost independent of the initial plasma conditions.
@@
The MHD model is evolved until the expected eruption of interest occurs, which takes approximately 3.2 hours. We set up a new MHD model to model the dynamics of the eruption.
II.2.3 Radiative MHD Model for the Eruption
@@
The third stage in the hybrid model is an MHD simulation with an extended vertical domain (hereafter, flare run). The flare run has a mesh of $512\times 512\times 3840$ grid points with the same grid spacing as the evo. run; thus, the domain height is extended to 245.76 Mm, which provides a large space for the upward eruption.
@@
The run cases discussed in this paper, including the production runs described above and control experiments that evaluate the effects of temporal and spatial scaling, are listed in Table 1.
[…]
Table 1: Summary of simulation cases
Case Name Grid Spacing Time Time Speed-up Volume
$N_{x}{\times}Nz$ $\Delta x$$\times\Delta z[km$] started evolved factor $f_{\rm sp}$ factor $f_{\rm V}$
zero-$\beta$ $256\times 1152$ $576\times 64$ Feb. 10 22h >4.5>4.5 days 12 1.8753
MHD evo. $512\times 1920$ $540\times 64$ Feb. 14 20h >3.2>3.2 hours 1 1
MHD flare $512\times 3840$ $540\times 64$ Feb. 14 23h >50>50 mins 1 1
[…]
Note. — All the models use $N_{x}=N_{y}$ and $\Delta x=\Delta y$.
III Results
III.1 Emergence of the Active Region and Energy Budget of the Erution
@@
The coronal magnetic field in the zero-$\beta$ run across 4 days [^2] The first three columns show the snapshots that is most close to $0^{\rm h}$ UT of each day, whereas the last column shows the snapshot corresponding to February 14, $20^{\rm h}$ UT, which is identical to the initial state of the MHD evo. simulation. are displayed in Panels (e) – (h) of Figure 1. The seed points for calculating fieldlines are uniformly distributed in the rectangular area covering the sunspots in this active region. In the latter two days, the coronal fieldlines closely resemble the shape of the observed EUV loops. However, a quantitative investigation of the formation of these loops requires a model with much finer resolution, as suggested in our earlier study on a non-eruptive active region (F. Chen, 2025).
A crucial parameter of flare-productive active regions is the free magnetic energy stored in the corona. Evolution of the volume-integrated total magnetic energy $E_{\rm tot}$, as given by
$E_{\rm tot}=\int\frac{1}{2\mu_{0}}\mathrm{B}^{2}\,dV,$ (4)
and free magnetic energy $E_{\rm free}$, which is estimated by
$E_{\rm free}=E_{\rm tot}-\int\frac{1}{2\mu_{0}}\mathrm{B^{2}_{\rm P}}\,dV,$ (5)
@@
in the zero-$\beta$ run is plotted in Panel (i) of Figure 1333On the basis of the setup of the zero-$\beta$ run, the magnetic energy and time are multiplied by a factor of $f_{\rm V}=1.875^{3}$ and $f_{\rm sp}=12$, respectively, such that they can be directly compared with quantities of the MHD simulations. $\mathbf{B}_{\rm P}$ stands for the potential field calculated from the vertical magnetic field at the bottom of the domain. The rapid increase in magnetic energy starts from February 13, similar to the trend found in observations (X. Sun et al., 2012). In our model, the magnetic energy increases to approximately $1.4\times 10^{33}$ erg with a very steady rate until a time corresponding to February 14, $23^{\rm h}$ UT; meanwhile, the free magnetic energy accumulates to approximately $6.5\times 10^{32}$ erg. We note that this energy includes a large contribution from the curved/twist magnetic field outside the core region of the domain, which is an effect of the constant $\Omega$ and the periodic side boundary. An abrupt decay in the free magnetic energy, as indicated by the gray mask in Panel (i), indicates a major eruption, which occurs approximately 3 hours earlier than the actual X class flare on real Sun (February 15 $2^{\rm h}$ UT). Given the 100-hour evolution time since February 10, 22h UT, this difference corresponds to an error of 3%.
@@
The evo. and flare runs describe the course of the eruption from development to onset. Figure 1(j) shows the joint evolution of the total and free magnetic energy since 2.5 hours before the flare onset. The time $t$ hereafter is relative to the beginning of the MHD flare run. The free magnetic energy steadily increases and starts to decrease at $t=219$ s. The fast decay lasts for approximately 1000 s, during which $1.94\times 10^{32}$ erg magnetic energy is released.
III.2 Magnetic Structure and Plasma Dynamics of the Eruption